# Relations Between the Inequality Indices Gini, Pietra and Kolkata: Theory and Data Analysis

### v1

@@ Acknowledgement
Categories Year Name of Nobel laureates Tot. no. of papers Tot. no. of citations Hirsch index ($h$) Gini index ($g$) Pietra index ($p$) Kolkata index ($k$) p/(2k1)p/(2k-1) $p/g$ (2k1)/g(2k-1)/g Economics 2020 Paul Milgrom 398 118799 83 0.9123 0.7965 0.8941 1.0105 0.8731 0.8640 2020 Robert Wilson 259 34825 58 0.8799 0.7379 0.8685 1.0012 0.8386 0.8376 2021 David Card 695 103999 117 0.8957 0.7627 0.8785 1.0075 0.8515 0.8452 2021 Guido Imbens 393 119042 102 0.8869 0.7522 0.8751 1.0025 0.8481 0.8459 2021 Josh Angrist 436 107740 92 0.9212 0.7891 0.8922 1.0060 0.8566 0.8515 2022 Ben Bernanke 717 147600 118 0.9210 0.7984 0.8984 1.0021 0.8669 0.8651 2022 Philip H. Dybvig 139 55939 41 0.9426 0.8339 0.9139 1.0073 0.8846 0.8782 2024 Daron Acemoglu 1358 277596 183 0.9167 0.7928 0.8939 1.0064 0.8649 0.8594 2024 James Robinson 845 131696 102 0.9465 0.8293 0.9146 1.0001 0.8761 0.8760 2024 Simon Johnson 947 95164 64 0.9693 0.8962 0.9477 1.0008 0.9246 0.9238 2025 Joel Mokyr 471 32139 71 0.9049 0.7548 0.8747 1.0071 0.8341 0.8282 2025 Philippe Aghion 647 146570 145 0.8733 0.7205 0.8586 1.0046 0.8250 0.8212 Physics 2020 Roger Penrose 639 101446 102 0.9022 0.7727 0.8859 1.0011 0.8565 0.8555 2021 Giorgio Parisi 1127 111103 136 0.8437 0.6694 0.8346 1.0002 0.7934 0.7933 2021 Syukuro Manabe 298 49325 89 0.8291 0.6687 0.8282 1.0186 0.8065 0.7918 2022 John F. Clauser 138 22474 33 0.9450 0.8298 0.9137 1.0029 0.8781 0.8756 2022 Aspect Alain 772 41611 76 0.9342 0.7984 0.8972 1.0050 0.8546 0.8503 2023 Anne LHuillier 514 36201 86 0.8470 0.6792 0.8375 1.0063 0.8018 0.7968 2023 Ferenc Krausz 1122 90527 130 0.8860 0.7335 0.8644 1.0064 0.8279 0.8226 2024 Geoffrey Hinton 751 975672 192 0.9440 0.8310 0.9155 1.0001 0.8803 0.8802 2024 John Hopfield 306 94961 95 0.8964 0.7381 0.8687 1.0009 0.8234 0.8227 2025 John Clarke 1081 50125 111 0.8501 0.6913 0.8409 1.0140 0.8132 0.8019 2025 Michel Devoret 754 78163 133 0.8586 0.7110 0.8478 1.0220 0.8281 0.8102 Chemistry 2020 Emmanuelle Charpentier 269 64602 65 0.9309 0.8030 0.9014 1.0002 0.8626 0.8624 2020 Jennifer Doudna 864 156101 164 0.8582 0.6868 0.8429 1.0014 0.8003 0.7992 2021 Benjamin List 337 49436 103 0.7482 0.5753 0.7868 1.0029 0.7690 0.7667 2021 David MacMillan 578 87442 134 0.8288 0.6704 0.8296 1.0171 0.8089 0.7952 2022 Carolyn Bertozzi 1021 103476 153 0.8081 0.6306 0.8132 1.0067 0.7804 0.7752 2022 Morten Meldal 412 32092 68 0.8261 0.6382 0.8165 1.0083 0.7726 0.7663 2023 Louis Brus 468 92751 122 0.8175 0.6606 0.8274 1.0088 0.8081 0.8010 2023 Moungi G. Bawendi 989 178538 189 0.8323 0.6666 0.8314 1.0057 0.8009 0.7964 2024 Demis Hassabis 171 238294 100 0.8530 0.6991 0.8492 1.0009 0.8196 0.8189 2024 David Baker 2690 201046 227 0.8431 0.6868 0.8399 1.0102 0.8146 0.8064 2025 Omar M. Yaghi 741 275500 197 0.8478 0.6895 0.8427 1.0061 0.8133 0.8084 2025 Susumu Kitagawa 1129 104460 146 0.7994 0.6226 0.8112 1.0002 0.7788 0.7787 Medicine 2020 Michael Houghton 546 61896 107 0.8429 0.6749 0.8366 1.0025 0.8007 0.7987 2022 Svante Paabo 587 151556 179 0.7805 0.6006 0.7989 1.0046 0.7696 0.7660 2023 Katalin Karikó 226 32576 66 0.8429 0.6874 0.8408 1.0084 0.8155 0.8087 2024 Victor Ambros 179 73309 72 0.8827 0.7315 0.8653 1.0011 0.8287 0.8278 2025 Shimon Sakaguchi 676 143255 136 0.8794 0.7323 0.8638 1.0065 0.8327 0.8274

### v2

@@ IV Summary and Conclusion
Categories
Year
Name of Nobel laureates
Tot. no. of papers
Tot. no. of citations
Hirsch index ($h$)
Gini index ($g$)
Pietra index ($p$)
Kolkata index ($k$)
p/(2k1)p/(2k-1)
$p/g$
(2k1)/g(2k-1)/g
Economics 2020 Paul Milgrom 398 118799 83 0.9123 0.7965 0.8941 1.0105 0.8731 0.8640
2020 Robert Wilson 259 34825 58 0.8799 0.7379 0.8685 1.0012 0.8386 0.8376
2021 David Card 695 103999 117 0.8957 0.7627 0.8785 1.0075 0.8515 0.8452
2021 Guido Imbens 393 119042 102 0.8869 0.7522 0.8751 1.0025 0.8481 0.8459
2021 Josh Angrist 436 107740 92 0.9212 0.7891 0.8922 1.0060 0.8566 0.8515
2022 Ben Bernanke 717 147600 118 0.9210 0.7984 0.8984 1.0021 0.8669 0.8651
2022 Philip H. Dybvig 139 55939 41 0.9426 0.8339 0.9139 1.0073 0.8846 0.8782
2024 Daron Acemoglu 1358 277596 183 0.9167 0.7928 0.8939 1.0064 0.8649 0.8594
2024 James Robinson 845 131696 102 0.9465 0.8293 0.9146 1.0001 0.8761 0.8760
2024 Simon Johnson 947 95164 64 0.9693 0.8962 0.9477 1.0008 0.9246 0.9238
2025 Joel Mokyr 471 32139 71 0.9049 0.7548 0.8747 1.0071 0.8341 0.8282
2025 Philippe Aghion 647 146570 145 0.8733 0.7205 0.8586 1.0046 0.8250 0.8212
Physics 2020 Roger Penrose 639 101446 102 0.9022 0.7727 0.8859 1.0011 0.8565 0.8555
2021 Giorgio Parisi 1127 111103 136 0.8437 0.6694 0.8346 1.0002 0.7934 0.7933
2021 Syukuro Manabe 298 49325 89 0.8291 0.6687 0.8282 1.0186 0.8065 0.7918
2022 John F. Clauser 138 22474 33 0.9450 0.8298 0.9137 1.0029 0.8781 0.8756
2022 Aspect Alain 772 41611 76 0.9342 0.7984 0.8972 1.0050 0.8546 0.8503
2023 Anne LHuillier 514 36201 86 0.8470 0.6792 0.8375 1.0063 0.8018 0.7968
2023 Ferenc Krausz 1122 90527 130 0.8860 0.7335 0.8644 1.0064 0.8279 0.8226
2024 Geoffrey Hinton 751 975672 192 0.9440 0.8310 0.9155 1.0001 0.8803 0.8802
2024 John Hopfield 306 94961 95 0.8964 0.7381 0.8687 1.0009 0.8234 0.8227
2025 John Clarke 1081 50125 111 0.8501 0.6913 0.8409 1.0140 0.8132 0.8019
2025 Michel Devoret 754 78163 133 0.8586 0.7110 0.8478 1.0220 0.8281 0.8102
Chemistry 2020 Emmanuelle Charpentier 269 64602 65 0.9309 0.8030 0.9014 1.0002 0.8626 0.8624
2020 Jennifer Doudna 864 156101 164 0.8582 0.6868 0.8429 1.0014 0.8003 0.7992
2021 Benjamin List 337 49436 103 0.7482 0.5753 0.7868 1.0029 0.7690 0.7667
2021 David MacMillan 578 87442 134 0.8288 0.6704 0.8296 1.0171 0.8089 0.7952
2022 Carolyn Bertozzi 1021 103476 153 0.8081 0.6306 0.8132 1.0067 0.7804 0.7752
2022 Morten Meldal 412 32092 68 0.8261 0.6382 0.8165 1.0083 0.7726 0.7663
2023 Louis Brus 468 92751 122 0.8175 0.6606 0.8274 1.0088 0.8081 0.8010
2023 Moungi G. Bawendi 989 178538 189 0.8323 0.6666 0.8314 1.0057 0.8009 0.7964
2024 Demis Hassabis 171 238294 100 0.8530 0.6991 0.8492 1.0009 0.8196 0.8189
2024 David Baker 2690 201046 227 0.8431 0.6868 0.8399 1.0102 0.8146 0.8064
2025 Omar M. Yaghi 741 275500 197 0.8478 0.6895 0.8427 1.0061 0.8133 0.8084
2025 Susumu Kitagawa 1129 104460 146 0.7994 0.6226 0.8112 1.0002 0.7788 0.7787
Medicine 2020 Michael Houghton 546 61896 107 0.8429 0.6749 0.8366 1.0025 0.8007 0.7987
2022 Svante Paabo 587 151556 179 0.7805 0.6006 0.7989 1.0046 0.7696 0.7660
2023 Katalin Karikó 226 32576 66 0.8429 0.6874 0.8408 1.0084 0.8155 0.8087
2024 Victor Ambros 179 73309 72 0.8827 0.7315 0.8653 1.0011 0.8287 0.8278
2025 Shimon Sakaguchi 676 143255 136 0.8794 0.7323 0.8638 1.0065 0.8327 0.8274